Find the inverse, g−1(t), when g(t) is given by 1/5t+10.

g−1(t)=
1
Expert's answer
2020-12-15T19:12:14-0500

Solution:

"g(t)= \dfrac{1}{5t+10}"


let`s call g(t) as y:


"y= \dfrac{1}{5t+10}" then "( \\dfrac{1}{y})\t-10=5t"


"t = \dfrac{1}{5y}-2"


Answer:

"g(t)^{-1} = \dfrac{1}{5t}-2" is inverse.